# CAMeleon: Interactively Exploring Craft Workflows in CAD

### v1

@@ 9. CONCLUSION AND FUTURE WORK
We have presented CAMeleon an interface to explore and compare workflows for computational fabrication. We furthermore introduce the idea of human-centered CAM workflows as synergistic workflows situated between classical CAM and manual craft processes. We have evaluated the usability and utility of CAMeleon with 12 participants in our study, demonstrating a variety of workflows on volumetric models they generated. We validated the ability to extend CAMeleon with new workflows, by collaborating with six craftspeople and implementing their workflows as human-centered CAM workflows in CAMeleon. We see CAMeleon as a small step towards a future where fabrication users share and explore workflows and processes, instead of specific models for fabrication. Similar to how recipe books for cooking serve as a starting point for inspiration instead of an endpoint to replicate the exact same dish.
Moving forwards we plan to further improve the extensibility of CAMeleon, now we have identified a series of commonly reusable modules and workflows, it would be interesting to create a visual block-based programming environment where craftspeople without coding experience can remix new workflows. We furthermore plan to distribute CAMeleon in local fabrication and craft communities to study it’s longer-term usage and adoption. And following recommendations from the study, integrating AI to generate models and browse workflows within CAMeleon could streamline the overall experience.
@@ Acknowledgements.
- Brooke (1996) john Brooke. 1996. SUS: A ’Quick and Dirty’ Usability Scale. In Usability Evaluation In Industry. CRC Press. Num Pages: 6.

### v2

@@ 11. CONCLUSION AND FUTURE WORK
We have presented CAMeleon a tool to explore and compare workflows for computational fabrication. We introduce the idea of human-centered CAM workflows as synergistic workflows between classical CAM and manual craft processes. In a workshop evaluation, we demonstrated CAMeleon’s utility and found that designing and exploring on a workflow level profoundly impacts the design process, and that it benefits designers at different levels of expertise. The modular architecture of CAMeleon supported our efforts to extend it with workflows from craftspeople and research projects evidences the extensibility of CAMeleon and helped identify future opportunities for technical implementation.
We see CAMeleon as a small step towards a future where fabrication users share and explore workflows and processes, instead of overly specific models for fabrication. This is similar to how recipe books for cooking serve as a starting point for inspiration instead of an endpoint to replicate the exact same dish.
Moving forward, we plan to advance the extensibility of CAMeleon by creating a visual programming environment where craftspeople without coding experience can remix new workflows and design their own. We furthermore plan to distribute CAMeleon in local fabrication and craft communities to study its longer-term usage and adoption.
@@ Acknowledgements.
- Ingold (2010) Tim Ingold. 2010. The textility of making. Cambridge journal of economics 34, 1 (2010), 91–102.
@@
- Polanyi (1967) M. Polanyi. 1967. The Tacit Dimension. Anchor Books. https://books.google.com/books?id=jwLXAAAAMAAJ
@@
- Sennett (2008) Richard Sennett. 2008. The craftsman. Yale University Press.
@@
- Zoran and Buechley (2013) Amit Zoran and Leah Buechley. 2013. Hybrid reassemblage: an exploration of craft, digital fabrication and artifact uniqueness. Leonardo 46, 1 (2013), 4–10.
@@ Appendix A Workflows Sources
Appendix B Step-by-Step Needle Felting Human-centered CAM Workflow